Find the total capacitance for three capacitors connected in series, given their individual capacitances
are
1
Expert's answer
2020-05-14T09:27:26-0400

Assume that we have three capacitors with corresponding capacitances "C_1,C_2, C_3". Let us find the total (equivalent) capacitance "C" of the series combination of our capacitors:


"\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3},\\
\space\\
C=\frac{C_1C_2C_3}{C_1C_2+C_2C_3+C_1C_3}."